A helium nucleus and neon nucleus are separated by a distance of 3.0 nanometers. The charge of the helium nucleus is +2e while that of neon nucleus is +10e. Find the magnitude of the electrostatic force they exert to one another.


1
Expert's answer
2022-03-22T15:47:14-0400

Electric force

"F=\frac{kq_1q_2}{r^2}"


"F=\frac{9\times10^9\times10\times1.6\times10^{-19}\times2\times1.6\times10^{-19}}{(3\times10^{-9})^2}=5.12\times10^{-10}N"